You are evaluating a project for The Tiff-any golf club, guaranteed to correct that nasty slice. You estimate the sales price of The Tiff-any to be $430 per unit and sales volume to be 1,200 units in year 1; 1,325 units in year 2; and 1,000 units in year 3. The project has a 3-year life. Variable costs amount to $240 per unit and fixed costs are $100,000 per year. The project requires an initial investment of $156,000 in assets, which will be depreciated straight-line to zero over the 3-year project life. The actual market value of these assets at the end of year 3 is expected to be $32,000. NWC requirements at the beginning of each year will be approximately 20 percent of the projected sales during the coming year. The tax rate is 30 percent and the required return on the project is 11 percent.
Answer #1

Sales price per unit of the product = $430 p.u.

Variable Cost per unit of the product = $240 p.u.

Contribution p.u of the product = Sales price per unit - Variable Cost per unit

= 430 – 240

= $190 p.u.

.

Fixed Cost per year = $100,000 per year

.

Initial Investment in Assets = 156,000

Depreciation every year = 156,000/3 = $52,000 every year for 3 years

Net Working Capital required & Inflow and Outflow from Working Capital (calculated in Table below):

Year Sales Units (1) Sales price p.u. (2) Total Sales Value ($) = (1)*(2) Working Capital required at the beginning of each year Inflow/Outflow of Working Capital
0                                     103,200              (103,200)
1 1200 430                                              516,000                                     113,950                 (10,750)
2 1325 430                                              569,750                                       86,000                   27,950
3 1000 430                                              430,000                                                -                     86,000

After Tax Salvage value on sale of asset = Profit on sale of Asset*(1-Tax Rate)

= (32,000-0)*(1-0.30)

= $22,400

.

Calculation of NPV:

Year Initial Investment (1) Contribution p.u. (2) Sales Units (3) Total Contribution (4) = (2)*(3) Fixed Cost (5) Depreciation(6) Profit Before Tax(7) = (4)-(5)-(6) Tax (8)= (7)*30% Profit after Tax (9) = (7)-(8) Cash Flow from Opeations(10) = (9)+(6) Working Capital (11) Salavge Value (12) Net Cash Flow (13) = (1) + (10) + (11)+(12) DF Working Discounting Factor (14) Present Value(15) = (13)*(14)
0 -156000                         (103,200)                                                                    (259,200) 1 1                                            (259,200)
1 0 190                   1,200                                                 228,000                                                 100,000                    52,000                                                      76,000                            22,800                                           53,200                                                                 105,200                            (10,750)                                                                         94,450 1/1.11^1 0.900900901                                                85,090
2 0 190                   1,325                                                 251,750                                                 100,000                    52,000                                                      99,750                            29,925                                           69,825                                                                 121,825                              27,950                                                                      149,775 1/1.11^2 0.811622433                                              121,561
3 0 190                   1,000                                                 190,000                                                 100,000                    52,000                                                      38,000                            11,400                                           26,600                                                                   78,600                              86,000                          22,400                                                                      187,000 1/1.11^3 0.731191381                                              136,733
NPV:                                                84,184

Company should invest in the project since NPV is positive .Project is viable.